OpenStudy (anonymous):

\[f(2x-3)=4x+1 \quad , \quad f(x+2)=?\] Attempted: I substituted 2x-3=t so... 2x-3=t \[x=\frac{t+3}{2}\] Now \[f(t)=4\cdot \left(\frac{t+3}{2}\right)+1\] f(t)=2t+7 now substitute t=x f(x)=2x+7 but how to find f(x+2)=? ..I need a hind please.

OpenStudy (anonymous):

i think you did most of the work already... since f(x) = 2x + 7, then f(x+2) = 2(x+2) + 7... isn't it?

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

... thank you, anyway i prefer to show my work , so the people can help me easily. There I was confused, I didn't know if I have to add +2 like this: f(x+2)=(2x+7)+2 or as you did... thanks again. :D

OpenStudy (anonymous):

I know the result I'll be the same ... but I thought about other cases... in General ;)